# Proposals for experimentally realizing (mostly) quantum-autonomous gates

### v1

@@ Abstract
Autonomous quantum machines (AQMs) execute tasks without requiring time-dependent external control. Motivations for AQMs include the restrictions imposed by classical control on quantum machines’ coherence times and geometries. Most AQM work is theoretical and abstract; yet an experiment recently demonstrated AQMs’ usefulness in qubit reset, crucial to quantum computing. To further reduce quantum computing’s classical control, we propose realizations of (fully and partially) quantum-autonomous gates on three platforms: Rydberg atoms, trapped ions, and superconducting qubits. First, we show that a Rydberg-blockade interaction or an ultrafast transition can quantum-autonomously effect entangling gates on Rydberg atoms. One can perform Z or entangling gates on trapped ions mostly quantum-autonomously, by sculpting a linear Paul trap or leveraging a ring trap. Passive lasers control these gates, as well as the Rydberg-atom gates, quantum-autonomously. Finally, circuit quantum electrodynamics can enable quantum-autonomous Z and XY gates on superconducting qubits. The gates can serve as building blocks for (fully or partially) quantum-autonomous circuits, which may reduce classical-control burdens.
@@ I Introduction
Autonomous quantum computation has attracted interest recently. In one approach (dissipative quantum computation), a quantum system couples to a tailored environment [27, 28, 29, 30, 31]. The environment guides the system toward a steady state that encodes a computation’s output. In an abstract-theory approach, researchers designed an autonomous quantum processing unit [32]. A quantum state encodes a program, which directs the processing unit to enact unitaries on the computational register. In another abstract-theory approach, Woods showed that quantum control can increase quantum gates’ speeds (as functions of resources such as energy) [33]. We complement these approaches, proposing experimental implementations of (mostly) quantum-autonomous gates.
We design mostly quantum-autonomous gates for three platforms: Rydberg atoms, trapped ions, and superconducting qubits. Section II shows how to perform quantum-autonomous entangling gates on Rydberg atoms, leveraging Rydberg-blockade interactions and ultrafast transitions. Section III details implementations of $Z$ and entangling gates on trapped ions. These mostly quantum-autonomous schemes rely on certain trapping potentials (sculpted linear Paul traps and ring Paul traps). Section IV describes quantum-autonomous $Z$ and $XY$ gates for superconducting qubits. These gates benefit from circuit quantum electrodynamics (QED). Section V outlines directions for future work. Our gates can serve as building blocks for (fully or partially) quantum-autonomous circuits.
@@ III Trapped ions
Here, we propose partially autonomous quantum gates for trapped ions. In Sec. III A, we overview conventional $Z$ and Mølmer–Sørensen (entangling) gates for trapped ions. We review ion traps, and compare them with the requirements of quantum autonomy, in Sec. III B. Section III C shows how tailored trapping potentials can enable partially quantum-autonomous $Z$ and Mølmer–Sørensen gates. Ring traps, too, can implement partially quantum-autonomous Mølmer–Sørensen gates, we show in the same subsection.
@@ III A Background: Trapped-ion quantum computing
In trapped-ion quantum computing, an ion’s internal (electronic) states serve as a qubit’s logical states. Species used include 40Ca+ [57], 137Ba+ [58], 171Yb+ [59], 88Sr+ [60], and 9Be+ [61]. A calcium ion’s $\lvert S_{1/2}\rangle$ and $\lvert D_{1/2}\rangle$ form a qubit, for example.
@@
III A 1 $Z$ gate
In reviewing the trapped-ion $Z$ gate, we follow [62]. [^8] An alternative scheme effects fast gates, which require times much shorter than the ion motion’s characteristic time scale [44, 63]. Our quantum-autonomous gates rely on ion-trap features suited to longer interaction times. We therefore focus on standard single-qubit gates, which last for microseconds. Let $\Omega$ denote an ion transition’s Rabi frequency. A $Z$ gate requires a laser whose frequency is detuned from $\Omega$ by some amount $\Delta\neq 0$. A laser wavelength of 397 nm serves. The laser shifts the ion’s transition frequency by $\delta=\frac{\Omega^{2}}{2\Delta}$. The shift leads to the unitary $e^{-i\theta\sigma_{z}/2}\,.$
@@ III A 2 Mølmer–Sørensen gate
This section reviews the Mølmer–Sørensen gate, which entangles two ions by coupling their internal states to a shared vibrational mode [64, 65]. Below, we detail the protocol and specify the laser resources required.
@@ III B Ion traps and quantum autonomy
In this subsection, we describe ion traps and their partial enabling of quantum autonomy. We focus on Paul traps, due to their commonness in quantum computing and for simplicity. However, our second gate proposal (Sec. III C 2) involves a ring trap, which can be a Penning trap [67].
Paul traps are the most popular ion traps applied in quantum computing. They confine ions using oscillating radio-frequency electric fields [67]. Paul traps shape ions into crystals whose geometries depend on the electrodes’ arrangement. Four cylindrical electrodes can form a linear Paul trap. Alternatively, the electrodes can lie in a plane, trapping ions above it [68]. By using more planar electrodes, one can form a microfabricated surface trap that generates an arbitrary potential. Examples include rings [69].
A Paul trap’s electric fields oscillate because the trap voltages vary. This variation changes the ions’ microscopic Hamiltonian; the classical trap performs thermodynamic work on the ions’ information-bearing DOFs. This changing and this work prevent the ions from acting as an AQM: by definition, the microscopic Hamiltonian responsible for undertaking an AQM’s mission must remain constant [2]. No classical control system may directly facilitate that mission by performing work. One might expect the trap merely to stabilize the ions’ spatial DOFs, rather than enacting gates. Such control could enable the ions to form an AQM. (Such control facilitates our Rydberg-atom and superconducting-qubit protocols, as discussed in App. B.) However, traps help enact gates in the next subsection’s protocols. For example, a ring trap spins ions in a circle, passing them through lasers that effect gates. Nevertheless, the lasers in our trapped-ion proposals—the tools that effect gates most directly—are quantum-autonomous, the next section shows. Hence the following gate protocols are partially quantum-autonomous.
III C Partially quantum-autonomous trapped-ion gates
Section III C 1 shows how to implement $Z$ and Mølmer–Sørensen gates partially quantum-autonomously, by tailoring a potential to form a slide. Ring trips can enable partially quantum-autonomous Mølmer–Sørensen gates, too, we show in Sec. III C 2.
@@ III C 1 Tailored potential slide
The proposal features the following setup. The trap forms a spatial gradient along which ions slide (Fig. 2). Laser beams, emitted by flash-lamp-pumped lasers, shine on particular locations. If an ion passes through a beam alone, it undergoes a single-qubit gate. If two ions pass through a beam simultaneously, they undergo a two-qubit gate.
Figure 2: Ions sliding down a potential gradient. Laser beams (yellow) implement gates.
The proposal involves a spatially varying potential and so an electric-field gradient. One can effect such a potential (or any other potential) using a Paul trap with an intricate-enough electrode structure [70]. Surface traps with multiple electrodes can serve this purpose [71, 72]. [^9] Our scheme differs from the ion-transport protocols in [73, 74, 75, 76, 77]. There, one adjusts direct-current voltages midprotocol. The adjustment shifts the trapping potentials’ minima, displacing the ions. This adjustment renders the protocols nonautonomous. Furthermore, several of the papers focus on transporting, rather than performing gates on, ions [73, 74, 75, 76]. The potential ramp transports the ions axially. Lasers drive the gate by coupling internal electronic states to radial motional modes. Therefore, the axial transport does not directly affect the ions’ internal states.
The gate is quantum-autonomous, to within the considerations of Sec. III B. The setup (laser placement and pumping, electrode arrangement, and loading of ions into the trap) takes place prior to the gate protocol. Instantiating the potential amounts to “pressing go” on the AQM. The ions will begin rolling down the potential slide, as nothing is pinning them to the top. The laser beams shine throughout the protocol, as detailed below. Hence the ion-and-laser system evolves under a time-independent Hamiltonian, apart from experiencing the trap’s electric-field oscillations.
One final component of the setup needs detailing: the flash-lamp-pumped lasers that effect the gates. A flash lamp is a gas-filled tube that emits light when triggered by a pulse  [45]. [^10] The flash lamp is not the laser; rather, it pumps the laser: in flash-lamp pumping, one charges a capacitor. A trigger electrode discharges the capacitor, which energizes the lamp’s gas. The excited gas flashes, pumping the laser’s gain medium. Flash-lamp-pumped lasers emit pulses that typically last for microseconds or milliseconds [78]. Hence these lasers can provide the microsecond-long pulses required for trapped-ion gates. [^11] In contrast, PMLLs (the quantum-autonomous lasers of Sec. II B 1) emit picosecond- and nanosecond-scale pulses. Furthermore, one can charge a flash lamp’s capacitor before implementing a gate. Therefore, flash-lamp-pumped lasers can enable quantum-autonomous control.
@@
The slide can effect also a Mølmer–Sørensen gate. During it, two ions slide together. They pass through a bichromatic laser beam detuned from their qubit and shared-motional-mode frequencies. The shared mode transfers excitations between the ions’ electronic DOFs.
@@ III C 2 Ring-trap autonomous gate
Like a potential slide, a ring trap (Sec. III B and [83, 69, 84]) enables a partially quantum-autonomous gate. First, we sketch the gate protocol. Then, we review equipment specifications required in similar nonautonomous setups. These specifications, we argue, are feasible in our setup. For simplicity, we focus on Mølmer–Sørensen gates; but our strategy can extend to $Z$ gates.
Figure 3 illustrates the partially quantum-autonomous Mølmer–Sørensen gate. Two ions rotate along a ring. A bichromatic laser beam illuminates two segments of the ring. As one ion enters one segment, the other ion enters the other segment. The ions pass into and out of the segments several times, until the laser is discharged, as in Sec. III C 1.
Figure 3: Partially quantum-autonomous Mølmer–Sørensen gate implemented with a ring trap. The black disks represent ions. Their trajectory forms the dashed line. A bichromatic laser shines on two ring segments, driving internal transitions in the ions.
To identify the required laser specifications, we review relevant features of two experiments [66, 84]. Both featured calcium ions. In one, the ions underwent a Mølmer–Sørensen gate [66]. The gate relied on a laser pulse that lasted for $\approx 50$ $\mu\text{s}$. The second experiment featured a ring trap [84]. Two ions rotated at a frequency of $100$ kHz on a circle of radius $r=$ 3.13 $\mu\text{m}$. [^13] The authors also achieved control crucial for two-qubit gates, indirectly supporting our proposal. One can entangle ions if they share a motional mode (Sec. III A). This mode should correspond to the sidebands addressed by the laser pulses. In a ring trap, the relevant degrees of freedom have a nonlinear energy spectrum, which complicates mode selection. Yet the authors addressed separated groups of spectral lines individually, by exciting the ions to a high-angular-momentum state. Hence the authors controlled the motional modes coherently, as is necessary for implementing gates.
Applying these numbers, we can begin to estimate, and argue for the practicality of, the equipment specifications required for our gate. First, the gate requires a laser pulse of $\approx 50$ $\mu\text{s}$ [66]. Flash-lamp-pumped lasers can produce pulses of this duration; their pulse lengths range from the microsecond to the millisecond scale [78].
Second, we estimate how many times the ions pass through the laser beams during this time interval. To do so, we must estimate the angle subtended by the laser beam (at one of its intersections with the ring) (Fig. 3). Reference [84] features a ring trap of radius $r\approx 3~\mu\text{m}$. Furthermore, beams with micron-scale diameters have been applied to trapped ions [79]. We therefore call for a beam diameter of a few microns, e.g., $\approx 4$ $\mu\text{m}$. This beam diameter, with the ring-trap radius, points to an angle $\theta\approx\pi/2$ subtended by the laser beam (Fig. 4).
[…]
Figure 4: Bird’s-eye view of ring trap and laser implementing a partially quantum-autonomous Mølmer–Sørensen gate. The beam diameter extends over a distance $l=2r\sin(\theta/2)$. Two ions simultaneously enter arcs illuminated by the laser beam.
[…]
@@ V Outlook
We have proposed implementations of (partially and fully) autonomous quantum gates for three experimental platforms: Rydberg atoms, trapped ions, and superconducting qubits. These gates could constitute building blocks for (partially or fully) autonomous quantum circuits. Greater autonomy could improve quantum devices’ scalability, coherence times, and energy costs.
@@ Acknowledgements.
The authors thank Joseph Britton, Marko Cetina, Andrew Goffin, Marcus Huber, Mikhail Lukin, and Matteo Mitrano for valuable conversations. J. A. M. G. acknowledges support from NIST grant 70NANB21H055_0. Y.-X. W. acknowledges support from a QuICS Hartree Postdoctoral Fellowship. T. M. acknowledges support from the Harvard Quantum Initiative Postdoctoral Fellowship in Science and Engineering. P. E. acknowledges funding from the Austrian Federal Ministry of Education, Science and Research via the Austrian Research Promotion Agency (FFG) through the project FO999914033 (QUICHE) and from the European Research Council (Consolidator grant “Cocoquest” 101043705). S. G. acknowledges financial support from the European Research Council (Grant No. 101041744 ESQuAT) and from the Knut and Alice Wallenberg Foundation through the Wallenberg Centre for Quantum Technology (WACQT). P. E. and S. G. are co-funded by the European Union (Quantum Flagship project ASPECTS, Grant Agreement No. 101080167). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union, REA, or UKRI. Neither the European Union nor UKRI can beheld responsible for them. N. L. and N. Y. H. acknowledge support from the National Science Foundation (QLCI grant OMA-2120757).
@@ Appendix B Time-dependent classical control in Rydberg-atom and superconducting-qubit proposals
Section III B detailed a subtlety of our trapped-ion proposals: Paul traps exert time-dependent classical control on the ions. The other platforms discussed (Rydberg atoms and superconducting qubits) require time-dependent classical control, too. Yet these two platforms’ control does not manipulate quantum-information-bearing DOFs, enacting gates. Hence our Rydberg-atom and superconducting-qubit gates qualify as quantum-autonomous, according to the definition in [2]. We support this claim for the two platforms consecutively.

### v2

# Proposals for experimentally realizing quantum-autonomous gates

@@ Abstract
Autonomous quantum machines (AQMs) execute tasks without requiring time-dependent external control. Motivations for AQMs include the restrictions imposed by classical control on quantum machines’ coherence times and geometries. Most AQM work is theoretical and abstract; yet an experiment recently demonstrated AQMs’ usefulness in qubit reset, crucial to quantum computing. To further reduce quantum computing’s classical control, we propose realizations of quantum-autonomous gates on three platforms: Rydberg atoms, trapped ions, and superconducting qubits. First, we show that a Rydberg-blockade interaction or an ultrafast transition can quantum-autonomously effect entangling gates on Rydberg atoms. Passive lasers control these gates quantum-autonomously. One can perform Z or entangling gates on trapped ions quantum-autonomously, by sculpting a linear Paul trap or leveraging a ring trap. Finally, circuit quantum electrodynamics can enable quantum-autonomous Z and XY gates on superconducting qubits. The gates can serve as building blocks for (fully or partially) quantum-autonomous circuits, which may reduce classical-control burdens.
@@ I Introduction
Autonomous quantum computation has attracted interest recently. In one approach (dissipative quantum computation), a quantum system couples to a tailored environment [27, 28, 29, 30, 31]. The environment guides the system toward a steady state that encodes a computation’s output. In an abstract-theory approach, researchers designed an autonomous quantum processing unit [32]. A quantum state encodes a program, which directs the processing unit to enact unitaries on the computational register. In another abstract-theory approach, Woods showed that quantum control can increase quantum gates’ speeds (as functions of resources such as energy) [33]. We complement these approaches, proposing experimental implementations of quantum-autonomous gates.
We design quantum-autonomous gates for three platforms: Rydberg atoms, trapped ions, and superconducting qubits. Section II shows how to perform quantum-autonomous entangling gates on Rydberg atoms, leveraging Rydberg-blockade interactions and ultrafast transitions. Section III details implementations of $Z$ and entangling gates on trapped ions. These quantum-autonomous schemes rely on certain trapping potentials (sculpted linear Paul traps and ring Paul traps). Section IV describes quantum-autonomous $Z$ and $XY$ gates for superconducting qubits. These gates benefit from circuit quantum electrodynamics (QED). Section V outlines directions for future work. Our gates can serve as building blocks for (fully or partially) quantum-autonomous circuits.
@@ III Trapped ions
Here, we propose autonomous quantum gates for trapped ions. In Sec. III A, we overview conventional $Z$ and Mølmer–Sørensen (entangling) gates for trapped ions. We review ion traps, and argue that they do not preclude quantum autonomy, in Sec. III B. Section III C shows how tailored trapping potentials can enable quantum-autonomous $Z$ and Mølmer–Sørensen gates. Ring traps, too, can implement quantum-autonomous $Z$ gates, as we show in the same subsection.
@@ III A Background: Trapped-ion quantum computing
In trapped-ion quantum computing, an ion’s internal (electronic) states serve as a qubit’s logical states. Species used include 40Ca+ [57], 137Ba+ [58], 171Yb+ [59], 88Sr+ [60], and 9Be+ [61]. For example, two states in a calcium ion’s $S_{1/2}$ and $D_{1/2}$ manifolds can form a qubit.
In the most common geometry, the ions form a chain. [^7] One of our proposals involves a ring geometry, but discussing chains suffices here. It has normal vibrational modes that can transfer information between different ions’ electronic degrees of freedom (DOFs). The modes correspond to a simple harmonic oscillator. Let $n$ denote the vibrational quantum number. We label the two ions’ electronic states by $\uparrow$ and $\downarrow$.
III A 1 $Z$ gate
@@ III A 2 Mølmer–Sørensen gate
This section reviews the Mølmer–Sørensen gate, which entangles two ions by coupling their internal states to a shared vibrational mode [64, 65]. Below, we overview the protocol and specify the laser resources required. We follow [65].
[…]
@@ III B Ion traps and quantum autonomy
In this subsection, we describe ion traps and their consistency with quantum autonomy. We focus on Paul traps, due to their commonness in quantum computing and for simplicity.
Paul traps are the most popular ion traps applied in quantum computing. They confine ions using oscillating radio-frequency, as well as static, electric fields [67]. Paul traps shape ions into crystals whose geometries depend on the electrode arrangement. Four cylindrical electrodes can form a linear Paul trap. Alternatively, the electrodes can lie in a plane, trapping ions above it [68]. By using more planar electrodes, one can form a microfabricated surface trap that generates an arbitrary potential. Experimentalists have realized ring-shaped ion arrangements using such electrodes [69].
A Paul trap is a classical system whose electric fields oscillate to confine the ions. The oscillation drives a current, consuming and performing work. One might worry that this work prevents the platform from acting as an AQM: no classical control system may directly facilitate an AQM’s mission by performing work, which would change the microscopic AQM Hamiltonian responsible for undertaking that mission [24, 2]. Yet the trap does not perform work on that microscopic Hamiltonian. The trap confines the ions by inducing a harmonic-potential term in the ions’ Hamiltonian. That term remains constant in time (although it varies across space in the protocol of Sec. III C 1). Hence the trap does not change the ions’ microscopic Hamiltonian, performing no work on the ions.
Two objections suggest themselves. First, the trap consumes work to maintain the oscillator term in the ions’ Hamiltonian. Yet the term remains constant; so the trap does not perform work on the ions, according to a common quantum-thermodynamic definition of work [24]. [^9] The ion experiences an effective pseudopotential modeled by a time-independent Hamiltonian [70, Eq. (6)], although the oscillating trapping fields consume work. Similarly limited control facilitates our Rydberg-atom and superconducting-qubit protocols, as detailed in App. B. Second, as the ions oscillate, they push and pull the electrodes’ electrons. The trap must counteract this effect to keep the voltage constant. Hence the trap draws work to increase or reduce the current. Yet this increase or reduction does not alter the ions’ Hamiltonian.
In summary, the trap does not prevent the ion system from acting as an AQM. Furthermore, quantum-autonomous lasers drive the gates, we argue in Sec. III C 1. Hence the following gate protocols are quantum-autonomous, according to the definition in [2].
III C Quantum-autonomous trapped-ion gates
Section III C 1 shows how to implement $Z$ and Mølmer–Sørensen gates quantum-autonomously, by tailoring a potential to form a slide. Ring traps can enable quantum-autonomous $Z$ gates, too, we show in Sec. III C 2.
@@ III C 1 Tailored potential slide
The proposal features the following setup. Part of the trap forms a spatial gradient along which ions slide (Fig. 2). Later, the trap levels out. Laser beams shine on particular locations. If an ion passes through a beam alone, it undergoes a single-qubit gate. If two ions pass through a beam simultaneously, they undergo a two-qubit gate.
Figure 2: Tailored potential slide. A potential (dashed curve) confines the ions before the protocol, then is switched off. The ions initially slide down a potential gradient, then traverse the gate zones. Laser beams (yellow) implement gates.
The proposal involves a spatially varying potential and so an electric-field gradient. One can effect such a potential (or any other potential) using a Paul trap with an intricate-enough electrode structure [71]. Surface traps with multiple electrodes can serve this purpose [72, 73]. Experimentalists have designed traps that feature many electrodes [74], rendering our proposal’s requirements plausible. The potential ramp transports the ions axially. This transport does not affect the ions’ internal states: the internal states couple, rather, to radial motional modes via lasers during entangling gates or only to the laser light during Z gates. [^10] Gates have been realized via the shuttling of ions [75, 76, 77]. Active electrode control enabled this transport, in contrast with our quantum-autonomous proposal.
The gate is quantum-autonomous, consistently with Sec. III B. The setup (laser placement and pumping, electrode arrangement, and loading of ions into the trap) occurs before the gate protocol. The ions are confined initially. When the protocol starts, the potential is switched off. This step amounts to “pressing go” on the AQM. Without additional control, the ions roll down the potential slide (left-hand side of Fig. 2). The ions then move along the flat potential, their center of mass maintaining a constant velocity. Coulomb repulsion increases the interatomic distance. The laser beams, detailed next, are applied throughout the protocol.
@@
[…]
@@ III C 2 Ring-trap autonomous gate
Like a potential slide, a ring trap (Sec. III B and [80, 69, 81]) enables a quantum-autonomous $Z$ gate. First, we sketch the gate protocol. Then, we review specifications required in similar nonautonomous setups. These specifications, we argue, are feasible in our setup.
Figure 3 illustrates the quantum-autonomous gate. The ions traverse a ring. [^14] We picture multiple ions, despite describing a single-qubit gate for the convenience of our feasibility analysis. A laser beam illuminates two segments of the ring. The ions pass into and out of the illuminated segments several times, until the laser is discharged, as in Sec. III C 1.
Figure 3: Quantum-autonomous $Z$ gate implemented with a ring trap. The black disks represent ions. Their trajectory forms the dashed line. A laser shines on two ring segments, driving internal transitions in the ions.
The ions can revolve quickly enough to effectively experience one coherent pulse throughout the protocol. Consider a $^{43}\text{Ca}^{+}$ ion with qubit states $\lvert\downarrow\rangle=\lvert 4S_{1/2}^{F{=}4,M_{F}{=}0}\rangle$ and $\lvert\uparrow\rangle=\lvert 4S_{1/2}^{F{=}3,M_{F}{=}0}\rangle$, following [82]. The desired pulse induces a relative phase between the ion state’s $\lvert\downarrow\rangle$ and $\lvert\uparrow\rangle$ terms. Passing through an illuminated arc, the ion experiences a short pulse. Suppose the qubit decoheres much more slowly than the ion revolves. The short pulses jointly resemble one long pulse, increasingly shifting the qubit’s relative phase.
Now, we support the supposition that the qubit decoheres much more slowly than the ion revolves. Reference [81] reports on an experiment in which two calcium ions revolved at a frequency of 100 kHz. In contrast, a calcium qubit decoheres at a rate $0.167$ Hz [82]. As 0.167 Hz $\ll$ 100 kHz, the time scales meet the required condition.
@@ V Outlook
We have proposed implementations of autonomous quantum gates for three experimental platforms: Rydberg atoms, trapped ions, and superconducting qubits. These gates could constitute building blocks for (partially or fully) autonomous quantum circuits. Greater autonomy could improve quantum devices’ scalability, coherence times, and energy costs.
@@ Acknowledgements.
The authors thank Joseph Britton, Marko Cetina, Andrew Goffin, Marcus Huber, Mikhail Lukin, and Matteo Mitrano for valuable conversations. J. A. M. G. acknowledges support from NIST grant 70NANB21H055_0. Y.-X. W. acknowledges support from a QuICS Hartree Postdoctoral Fellowship. T. M. acknowledges support from the Harvard Quantum Initiative Postdoctoral Fellowship in Science and Engineering. P. E. acknowledges funding from the Austrian Federal Ministry of Education, Science and Research via the Austrian Research Promotion Agency (FFG) through the project FO999914033 (QUICHE) and from the European Research Council (Consolidator grant “Cocoquest” 101043705). S. G. acknowledges financial support from the European Research Council (Grant No. 101041744 ESQuAT) and from the Knut and Alice Wallenberg Foundation through the Wallenberg Centre for Quantum Technology (WACQT). P. E. and S. G. are co-funded by the European Union (Quantum Flagship project ASPECTS, Grant Agreement No. 101080167). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union, REA, or UKRI. Neither the European Union nor UKRI can beheld responsible for them. N. M. L. and N. Y. H. acknowledge support from the National Science Foundation (QLCI grant OMA-2120757).
@@ Appendix B Time-dependent classical control in Rydberg-atom and superconducting-qubit proposals
Section III B detailed a subtlety of our trapped-ion proposals: Paul traps are classical systems that consume and perform work, although not on the ions. The other platforms discussed (Rydberg atoms and superconducting qubits) require time-dependent classical control, too. Neither does these two platforms’ control manipulate quantum-information-bearing DOFs, enacting gates. Hence our Rydberg-atom and superconducting-qubit gates qualify as quantum-autonomous, according to the definition in [2]. We support this claim for the two platforms consecutively.